if -13 is added to a number and the sum is doubled the result is 14 less than the number. what is the number

Answer by ewatrrr(24785)


Hi,
Let x be the number
writing as reading the queston
2(x -13) = x -14
Simplify and Solve
2x - 26 = x - 14
x = 12
checking our answer
2*-1 = 12 -14 = -2